OpenStudy (anonymous):

Factor: \[1-\frac{ \sin^2x }{ 1+ cosx }\]

hartnn (hartnn):

write sin^2 x as 1-cos^2 x

hartnn (hartnn):

then use \( a^2-b^2=(a+b)(a-b)\)

hartnn (hartnn):

which gives you \(\huge \sin^2x=(1-\cos x)(1+\cos x)\)

hartnn (hartnn):

did u get that ?

hartnn (hartnn):

@johnnyalln a response is expected.....

OpenStudy (anonymous):

Got ya! Yes i was trying to write it all out as you explained

hartnn (hartnn):

ok...good, so what u gor finally ?

hartnn (hartnn):

*got

OpenStudy (anonymous):

still working it :) one sec

OpenStudy (anonymous):

I just get cos x...

hartnn (hartnn):

1-(1-cos x) = 1-1+cos x =cos x you are correct :)

OpenStudy (anonymous):

awesome! thank you so much!

hartnn (hartnn):

welcome :)